Four heater, each having a resistance of 30Ω, are connected in series across a 600v train circuit. 

Find the current?


1
Expert's answer
2021-09-16T17:04:38-0400

Gives


"R_1=R_2=R_3=R_4=R=30\Omega"

V=600volt


"R_{net}=R_1+R_2+R_3+R_4\\R_{net}=30+30+30+30=120\Omega"

Current

"I=\frac{v}{R_{net}}=\frac{300}{120}=2.5Amp"